OpenStudy (anonymous):

The following volume measurement are intial volume Final Volume 7.6 19.3 8.0 20.1 6.2 18.4 Comment on the accuracy of the volume measurements obtained by water displacement

OpenStudy (anonymous):

7.6/19.3 = 0.39 8.0/20.1 = 0.40 6.2/18.4 = 0.34 Not accurate i'd say. is that what you are supposed to do?

OpenStudy (anonymous):

what would be the percent accuracy for this measurments?

OpenStudy (ookawaiioo):

multiple by 100.....to get percentage

OpenStudy (anonymous):

Thanks guys!